Given (x^2+4x+8)divided by (x+k), find the values of k so that the remainder is 20.
Please use synthetic division.

Answer by lwsshak3(11628)

Given (x^2+4x+8)divided by (x+k), find the values of k so that the remainder is 20.
Please use synthetic division.
***
...-k..|....1........4.............8......
......................-k.......(-4k+k^2)
..............1....(4-k)....(8-4k+k^2)
8-4k+k^2=20
k^2-4k-12=0
(k-6)(k+2)=0
k=6
or
k=-2